A certain block is floating in water with 77% of itself submerged. It has length of 4.3 cm, a width of 5.9 cm, and a height of 2.75 cm. What is the mass, in kilograms?
1
Expert's answer
2020-08-31T07:05:02-0400

The buoyance force is equal to the weight of the object or the force of gravity, FB = Fg                  

Fg = VB*PB*g (force of gravity or the weight of the object = to volume of the object * density of the object * gravity)

FB = VS * Pw * g (the buoyant force = the volume of the submerged object * density of water * gravity)

Volume of the object = l*w*h = 4.3 cm * 5.9cm * 2.75cm = 69.7675cm3

77% volume of the submerged object = 77/100 * 69.7576 = 53.7210

Therefore, the weight of the object will be, 53.7210 cm3 * 1000 kg/m3 * 10 m/s2 = 0.5372 kg/m/s2

Mass of the object = 0.5372 kg/m/s2 divide by 10 m/s2     = 0.053721kg